# thanks: Authors of equal contributionsthanks: Authors of equal contributions

### v2

@@ Abstract
Non-Markovianity, the intricate dependence of an open quantum system on its temporal evolution history, holds tremendous implications across various scientific disciplines. However, accurately characterizing the complex non-Markovian effects has posed a formidable challenge for numerical simulations. Despite the promising potential of emerging quantum computing technologies, the pursuit of a universal theory enabling practical implementation of quantum computation algorithms remains ongoing. In this paper, we present a major advancement in bridging this critical gap: the development of dissipaton-embedded quantum master equation in second quantization (DQME-SQ). The DQME-SQ is an exact and compact theory, and its unique capabilities are demonstrated through digital quantum simulations of non-Markovian dissipative dynamics in both bosonic and fermionic environments. Our new theoretical developments pave the way for efficient exploration of complex open quantum systems.
Introduction.
@@
The simulation of open quantum systems (OQS) has attracted wide attention, because their interactions with the surrounding environment lead to intricate dynamical behavior Wei21 ; Bre02 ; Kle09 . OQS often exhibit non-Markovian features, characterized by memory effects where past states influence the current state. Understanding and characterizing non-Markovian effects are crucial for various physical processes such as quantum transport, quantum dephasing, coherent energy transfer, and precise measurement and control of quantum states Nit06 ; Bro16045005 ; Zhe13086601 ; Bre16021002 ; Dev17015001 . For example, in biomolecular systems such as the Fenna-Matthews-Olson complex and the B800-850 ring, the non-Markovianity facilitates highly efficient energy conversion and charge transfer Lee071462 ; Eng07782 ; Nal15022706 ; Kun22015101 . In strongly correlated quantum dots, non-Markovian memory leads to memristive behavior that may be harnessed in novel quantum device applications Zhe13086601 . However, characterizing non-Markovian effect in OQS has remained a longstanding challenge.
@@
Among the numerous attempts Tan906676 ; Tan20020901 ; Yan04216 ; Xu07031107 ; Sha045053 ; Sto02170407 ; Moi13134106 ; Han19050601 ; Tam18030402 ; Tam19090402 ; Lam193721 ; Sue14150403 ; Zha23080901 ; Ke22194102 ; Xu22230601 ; Ke16024101 to address this challenge, a representative method, the hierarchical equations of motion (HEOM) method Tan20020901 , has gained significant attention. The HEOM method characterizes non-Markovianity through a two-time correlation function in the form of a memory kernel, which is decomposed and reorganized within a hierarchy to account for all the possible dissipation modes. However, the exponential growth of dynamical variables presents serious challenges in computational time and data storage, making it difficult to apply HEOM and other non-Markovian quantum dissipation methods to complex OQS within existing classical computational frameworks Tan20020901 .
@@
In recent years, the rapid progress in quantum technologies has launched a new era of computational capabilities, enabling exploration of previously inaccessible realms of the modern science Van044 ; Lad107285 ; Bia17195 ; Cao1910856 ; Mca201 ; Oll20260511 ; Mie2325 ; Dal22667 ; Bla12277 ; Gro17995 . Quantum computation holds promise for efficiently simulating OQS. Several quantum computing algorithms have been developed for simulating quantum dynamics Oll20260511 ; Kam22010320 ; Mie2325 ; Cir23Arxiv2311_15240 ; Lam23Arxiv2310_12539 ; Zha24054101 ; Bar11486 ; Bar22043161 ; Li23147 ; Gac24013143 , including full quantum algorithms Sch21270503 ; Sch22023216 ; Hea21013182 and quantum-classical hybrid algorithms based on the variational principle Sar05250503 ; End20010501 ; Yao21030307 ; Wan234851 . However, simulating non-Markovian OQS in quantum circuits remains a challenging task. In the HEOM method, the main difficulty lies in the complex structure of the coupled equations, leading to intricate quantum circuits representing the propagator.
@@
This paper aims to establish a universal and compact theory, the dissipaton-embedded quantum master equation in second quantization (DQME-SQ), which enables the quantum simulation of non-Markovian open quantum dynamics.
@@
Construction of DQME-SQ theory.
@@
In our previous work Li23214110 , we had established the dissipaton-embedded quantum master equation (DQME). This theory effectively decomposes the non-Markovian environment into a set of statistically independent generalized Brownian quasi-particles, referred to as dissipatons Yan14054105 . Depicted in Fig. 1(a) is the total composite system, consisting of the system of primary interest and its non-Markovian environment. In Fig. 1(b), the environment is represented by a series of dissipatons (red plates). The system interacts with these dissipatons, with each dissipaton carrying non-Markovian memory information through its characteristic energy and lifetime. The dissipaton behaves like a Brownian particle embedded in a Markovian bath (light blue dots). The dynamics of the extended system, consisting of the system and dissipatons, is governed by the DQME, which is a time-local Fokker-Planck-like equation Li23214110 . Taking the bosonic open quantum system as the example, the temporal evolution of density distribution of the extended system, $\rho(x_{1},x_{2},\cdots)$, is governed by the DQME, where $x_{1},x_{2},\cdots$ are the coordinates of dissipatons. However, since the coordinates $\{x_{k}\}$ are continuous variables, it is difficult to encode the DQME for quantum simulation purposes.
@@
To overcome this difficulty, we notice that the quantum dissipation process can be fully characterized by a combination of generation and annihilation events of dissipatons. Such characterization corresponds to a second quantization procedure. In Fig. 1(c), each level corresponds to a dissipaton mode capable of accommodating multiple dissipatons. A dissipaton configuration is established by specifying the occupation numbers of all dissipaton modes. After applying the application of second quantization, the DQME-SQ governs the transitions between different Fock states of dissipatons. In principle, the non-Markovianity is exactly accounted for if all the possible dissipaton configurations are explicitly included in the open quantum dynamics. This is analogous to the electronic structure theory where full configuration interaction treatment yields exact many-body wavefunction.
@@
The construction of DQME-SQ theory will first be exemplified with bosonic open quantum system. We always set $\hbar=1$ and $\beta\equiv 1/(k_{\rm B}T)$, where $k_{\rm B}$ is the Boltzmann constant and $T$ the environment temperature. The total Hamiltonian has the generic form of $H_{\mbox{\tiny T}}=H_{\mbox{\tiny S}}+H_{\mbox{\tiny E}}+H_{\mbox{\tiny S}\mbox{\tiny E}}$. Here, $H_{\mbox{\tiny S}}$ and $H_{\mbox{\tiny E}}$ are the system and environment Hamiltonian, respectively. For brevity, the system–environment coupling is set as $H_{\mbox{\tiny S}\mbox{\tiny E}}=\hat{Q}\hat{F}$, where $\hat{Q}$ is the system operator, and the environment operator $\hat{F}$ incorporates the system–environment coupling strength. The non-Markovian memory of a Gaussian environment is fully characterized by the hybridization correlation function, $C(t)=\langle\hat{F}(t)\hat{F}(0)\rangle_{\mbox{\tiny E}}$. Here, $\hat{F}(t)\equiv e^{ih_{\mbox{\tiny E}}t}\hat{F}e^{-ih_{\mbox{\tiny E}}t}$ and $\langle\hat{O}\rangle_{\mbox{\tiny E}}\equiv{\rm Tr}_{\mbox{\tiny E}}[\hat{O}e^{-\beta H_{\mbox{\tiny E}}}]/{\rm Tr}_{\mbox{\tiny E}}(e^{-\beta H_{\mbox{\tiny E}}})$ is the ensemble average at the equilibrium state of bare environment, where $\hat{O}$ is an arbitrary operator.
@@
The DQME-SQ formalism starts with the dissipaton decomposition (DD) [cf. Arrow(1) in Fig. 1], which should preserve the correlation function $C(t)$. The DD maps the environment hybridization operator $\hat{F}$ onto $\{\hat{f}_{k}\}$, the coordinate operators of dissipatons Yan14054105 ,
@@
For individual dissipatons, we set $\langle\hat{f}_{k}(t)\hat{f}_{k^{\prime}}(0)\rangle_{\mbox{\tiny D}}=\delta_{kk^{\prime}}c_{k}(t)$, such that the non-Markovian memory of the original bosonic environment is fully recovered through the statistics of dissipatons, as
@@
Here, $\langle\hat{O}\rangle_{\mbox{\tiny D}}$ denotes the expectation value taken at the dissipaton vacuum state Wan22044102 . In this work, we adopt $c_{k}(t)=\eta_{k}e^{-\gamma_{k}t}$, lead to an exponential decomposition of $C(t)$, that can be accurately and routinely realized using the Prony decomposition algorithm. The resulting DD protocol is universal for arbitrary spectral density at any temperatures Che22221102 . The exponents are either real or complex conjugate paired Yan14054105 . Therefore, we may determine $\bar{k}$ by the pairwise equality $\gamma_{\bar{k}}=\gamma_{k}^{\ast}$. In the latter case, the dissipaton acquires both the characteristic energy and decaying rate, as the imaginary and real parts of $\gamma_{k}$, respectively. In the former case, $\gamma_{k}$ is real ($\bar{k}=k$) and the dissipaton mode purely decays.
The dissipaton is further second-quantized by the operator transformation
$\hat{f}_{k}\stackrel{{\scriptstyle\text{SQ}}}{{\longmapsto}}\zeta_{k}\,\big{(}\hat{b}_{k}^{-}+\hat{b}_{k}^{+}\big{)},$ (3)
where $\{\hat{b}_{k}^{+}\}$ and $\{\hat{b}_{k}^{-}\}$ are bosonic creation and annihilation operators satisfying $[\hat{b}^{-}_{k},\hat{b}^{+}_{k^{\prime}}]=\delta_{kk^{\prime}}$, with the choice of coefficient $\zeta_{k}=\sqrt{(\eta_{k}+\eta^{*}_{\bar{k}})/2}$. For each dissipaton, the Fock space ${\cal H}_{f_{k}}$ is then generated, and one can thus establish the DQME-SQ for the dynamical variable $\tilde{\rho}(t)$ of the extended system. Specifically, $\tilde{\rho}(t)$ is represented as a density tensor in the reduced system space ${\cal H}_{\mbox{\tiny S}}\otimes{\cal H}^{\ast}_{\mbox{\tiny S}}$ further dilated by the dissipaton Fock space ${\cal H}_{\mbox{\tiny D}}=\otimes_{k}{\cal H}_{f_{k}}$, i.e., $\tilde{\rho}\in{\cal H}_{\mbox{\tiny S}}\otimes{\cal H}^{\ast}_{\mbox{\tiny S}}\otimes{\cal H}_{\mbox{\tiny D}}$. The constructed DQME-SQ reads

### v3

@@ Abstract
Non-Markovianity, the intricate dependence of an open quantum system on its temporal evolution history, holds tremendous implications across various scientific disciplines. However, accurately characterizing the complex non-Markovian effects has posed a formidable challenge for numerical simulations. While quantum computing technologies show promise, a universal theory enabling practical quantum algorithm implementation has been elusive. We address this gap by introducing the dissipaton-embedded quantum master equation in second quantization (DQME-SQ). This exact and compact theory offers two key advantages: representability by quantum circuits and universal applicability to any Gaussian environment. We demonstrate these capabilities through digital quantum simulations of non-Markovian dissipative dynamics in both bosonic and fermionic environments. The DQME-SQ framework opens a new horizon for the efficient exploration of complex open quantum systems by leveraging the rapidly advancing quantum computing technologies.
I Introduction
@@
The simulation of open quantum systems (OQSs) has attracted wide attention, because their interactions with the surrounding environment lead to intricate dynamical behavior Wei21 ; Bre02 ; Kle09 . OQSs often exhibit non-Markovian features, characterized by memory effects where past states influence the current state. Understanding and characterizing non-Markovian effects are crucial for various physical processes such as quantum transport, quantum dephasing, coherent energy transfer, and precise measurement and control of quantum states Nit06 ; Bro16045005 ; Zhe13086601 ; Bre16021002 ; Dev17015001 . For example, in biomolecular systems such as the Fenna-Matthews-Olson complex and the B800-850 ring, the non-Markovianity facilitates highly efficient energy conversion and charge transfer Lee071462 ; Eng07782 ; Nal15022706 ; Kun22015101 . In strongly correlated quantum dots, non-Markovian memory leads to memristive behavior that may be harnessed in novel quantum device applications Zhe13086601 . However, accurate characterization of non-Markovian effect in complex OQSs has remained a longstanding challenge.
@@
Among the numerous attempts Sha045053 ; Sto02170407 ; Moi13134106 ; Han19050601 ; Gar974636 ; Tam18030402 ; Tam19090402 ; Lam193721 ; Sue14150403 ; Zha23080901 ; Ke22194102 ; Xu22230601 ; Ke16024101 ; Luo23030316 to address this challenge, a representative method, the hierarchical equations of motion (HEOM) method Tan89101 ; Tan906676 ; Tan20020901 ; Yan04216 ; Xu07031107 , has gained significant attention. The HEOM method characterizes non-Markovianity through a two-time correlation function in the form of a memory kernel, which is decomposed and reorganized within a hierarchy to account for all the possible dissipation modes. However, the exponential growth of dynamical variables presents serious challenges in computational time and data storage, making it difficult to apply HEOM and other non-Markovian quantum dissipation methods to complex OQSs within existing classical computational frameworks Tan20020901 .
@@
In recent years, the rapid progress in quantum technologies has launched a new era of computational capabilities, enabling exploration of previously inaccessible realms of the modern science Van044 ; Lad107285 ; Bia17195 ; Cao1910856 ; Mca201 ; Oll20260511 ; Mie2325 ; Dal22667 ; Bla12277 ; Gro17995 . Quantum computation holds promise for efficiently simulating OQSs. Several quantum computing algorithms have been developed for simulating quantum dynamics Oll20260511 ; Kam22010320 ; Mie2325 ; Cir23Arxiv2311_15240 ; Lam23Arxiv2310_12539 ; Zha24054101 ; Bar11486 ; Bar22043161 ; Li23147 ; Gac24013143 ; Sur231002 , including full quantum algorithms Sch21270503 ; Sch22023216 ; Hea21013182 and quantum-classical hybrid algorithms based on the variational principle Sar05250503 ; End20010501 ; Yao21030307 ; Wan234851 ; Guo24_arxiv_2404_10655 . However, simulating non-Markovian OQSs in quantum circuits remains a challenging task. In the HEOM method, the main difficulty lies in the complex structure of the coupled equations, leading to intricate quantum circuits representing the propagator. A novel formalism is required for the quantum simulation of non-Markovian open quantum dynamics.
@@
This work aims to establish a universal and compact theory, the dissipaton-embedded quantum master equation in second quantization (DQME-SQ). Its unique capabilities are demonstrated through digital quantum simulations of non-Markovian dissipative dynamics in both bosonic and fermionic environments.
The remainder of this paper is organized as follows. In Sec. II and Sec. III, we present the construction of DQME-SQ theories for bosonic and fermionic environments, respectively. The quantum simulation of non-Markoivan OQSs using DQME-SQ is elaborated in Sec. IV, with numerical demonstrations in Sec. V. Finally, we summarize the paper in Sec. VI.
@@
II Bosonic DQME-SQ theory
II.1 Construction of bosonic DQME-SQ
@@
In our previous work Li23214110 , we had established the dissipaton-embedded quantum master equation (DQME). In this theory, the non-Markovian memory content of the environment is effectively represented by a set of statistically independent generalized Brownian quasi-particles, referred to as dissipatons Yan14054105 . Depicted in Fig. 1(a) is the total composite system, consisting of the system of primary interest and its non-Markovian environment. In Fig. 1(b), the environment is represented by a series of dissipatons (red plates) that interact with the system. Each dissipaton behaves as a Brownian particle moving within a Markovian bath (light blue dots). The non-Markovian memory of the environment is manifested through the characteristic energies and lifetimes of the dissipatons. These dissipatons can interact strongly with the system, and are thus regarded as being embedded within the system. In the first quantization formulation, the DQME is a time-local Fokker-Planck-like equation Li23214110 . Taking the bosonic open quantum system as the example, the temporal evolution of density distribution of the dissipaton-embedded system, $\rho(x_{1},x_{2},\cdots)$, is governed by the DQME, where $x_{1},x_{2},\cdots$ are the coordinates of dissipatons. However, since the coordinates $\{x_{k}\}$ are continuous variables, it is difficult to encode the DQME for quantum simulation purposes.
To overcome this difficulty, we notice that the quantum dissipation process can be fully characterized by a combination of generation and annihilation events of dissipatons. Such characterization corresponds to a second quantization procedure. In Fig. 1(c), each short black bar corresponds to an energy level capable of accommodating any number of dissipatons, while the broadening of the level (represented by red shade) indicates the associated lifetime of dissipatons. A dissipaton configuration is thus established by specifying the occupancy on all the dissipaton levels.
@@
After applying the second quantization, the resulting DQME-SQ governs the transitions between different dissipaton configurations. In principle, the non-Markovianity is exactly accounted for if all the possible dissipaton configurations are explicitly included in the open quantum dynamics. This is analogous to the electronic structure theory where full configuration interaction treatment yields exact many-body wavefunction.
@@
The construction of DQME-SQ will first be exemplified with an OQS with a bosonic environment. In the following, we set $\hbar=1$ and $\beta\equiv 1/(k_{\rm B}T)$, where $k_{\rm B}$ is the Boltzmann constant and $T$ the environmental temperature. The total Hamiltonian has the generic form of $H_{\mbox{\tiny T}}=H_{\mbox{\tiny S}}+H_{\mbox{\tiny E}}+H_{\mbox{\tiny S}\mbox{\tiny E}}$. Here, $H_{\mbox{\tiny S}}$ and $H_{\mbox{\tiny E}}$ are the system and environment Hamiltonian, respectively. For brevity, the system–environment coupling is set as $H_{\mbox{\tiny S}\mbox{\tiny E}}=\hat{Q}\hat{F}$, where $\hat{Q}$ is a system operator, and the environment operator $\hat{F}$ incorporates the system–environment coupling strength. The non-Markovian memory of a Gaussian environment is fully characterized by the hybridization correlation function, $C(t)=\langle\hat{F}(t)\hat{F}(0)\rangle_{\mbox{\tiny E}}$. Here, $\hat{F}(t)\equiv e^{iH_{\mbox{\tiny E}}t}\hat{F}e^{-iH_{\mbox{\tiny E}}t}$, and $\langle\hat{O}\rangle_{\mbox{\tiny E}}\equiv{\rm Tr}_{\mbox{\tiny E}}(\hat{O}e^{-\beta H_{\mbox{\tiny E}}})/{\rm Tr}_{\mbox{\tiny E}}(e^{-\beta H_{\mbox{\tiny E}}})$ is the ensemble average at the equilibrium state of bare environment, where $\hat{O}$ is an arbitrary operator.
@@
The DQME-SQ formalism starts with the dissipaton decomposition (DD) [cf. Arrow (1) in Fig. 1], which should preserve the correlation function $C(t)$. The DD maps the environment hybridization operator $\hat{F}$ onto $\{\hat{f}_{k}\}$, the coordinate operators of dissipatons Yan14054105 ; Wan22170901 ,
@@
For individual dissipatons, we set $\langle\hat{f}_{k}(t)\hat{f}_{k^{\prime}}(0)\rangle_{\mbox{\tiny D}}=\delta_{kk^{\prime}}c_{k}(t)$, where $\langle\hat{O}\rangle_{\mbox{\tiny D}}$ denotes the expectation value taken at the dissipaton vacuum state Wan22044102 . It is important that the non-Markovian memory of the original bosonic environment is fully recovered through
@@
In this work, we adopt the function form of $c_{k}(t)=\eta_{k}e^{-\gamma_{k}t}$, which leads to an exponential decomposition of $C(t)$. Numerically, such a decomposition can be accurately and routinely realized using the Prony decomposition algorithm Che22221102 or other algorithms Xu22230601 ; Tak24204105 . The resulting DD protocol is universal for arbitrary spectral density of the environment and for any temperature Che22221102 . It has been proved that the exponents $\{\gamma_{k}\}$ are either real numbers or emerge as complex conjugate pairs, and the coefficients $\{\eta_{k}\}$ generally take complex values Yan14054105 ; Yan16110306 . Therefore, we may determine $\bar{k}$ by the pairwise equality $\gamma_{\bar{k}}=\gamma_{k}^{\ast}$, and the imaginary and real parts of $\gamma_{k}$ correspond to the characteristic energy and decaying rate of a dissipaton, respectively.
The coordinate operators of the dissipatons are further second-quantized by the transformation
$\hat{f}_{k}\stackrel{{\scriptstyle\text{SQ}}}{{\longmapsto}}\zeta_{k}\,\big{(}\hat{b}_{k}^{-}+\hat{b}_{k}^{+}\big{)}.$ (3)
Here, the coefficients are chosen as $\zeta_{k}=\sqrt{(\eta_{k}+\eta^{*}_{\bar{k}})/2}$. $\{\hat{b}_{k}^{+}\}$ and $\{\hat{b}_{k}^{-}\}$ are bosonic creation and annihilation operators, respectively. They satisfy the relation $[\hat{b}^{-}_{k},\hat{b}^{+}_{k^{\prime}}]=\delta_{kk^{\prime}}$. For each dissipaton, the Fock space ${\cal H}_{f_{k}}$ is then generated, and one can thus establish the DQME-SQ for the reduced density tensor (RDT) of the dissipaton-embedded system, denoted by $\tilde{\rho}(t)$. Here, $\tilde{\rho}$ spans the tensor product space of the original system and the dissipaton Fock space, i.e., $\tilde{\rho}\in{\cal H}_{\mbox{\tiny S}}\otimes{\cal H}^{\ast}_{\mbox{\tiny S}}\otimes{\cal H}_{\mbox{\tiny D}}$. The DQME-SQ for $\tilde{\rho}(t)$ reads
@@
The DQME-SQ of Eq. (II.1) yields exact non-Markovian dynamics for any Gaussian environment, provided that the hybridization correlation function of the environment can be decomposed into a sum of exponential functions. This exactness will be validated in Sec. II.2 by establishing a formal equivalence between Eq. (II.1) and the bosonic HEOM. Moreover, it will also be demonstrated in Sec. II.3 that Eq. (II.1) is universally applicable to general OQSs, where the system contains multiple degrees of freedom and is coupled to more than one environment.
The pseudomode theory Gar974636 ; Tam18030402 ; Tam19090402 ; Lam193721 ; Cir23Arxiv2311_15240 ; Lam23Arxiv2310_12539 has been established based on an exponential decomposition of the environment correlation function $C(t)$, leading to a Lindblad equation capable of capturing the exact non-Markovian open quantum dynamics, subject to the condition that the pre-exponential coefficients $\{\eta_{k}\}$ are real numbers, as implied by Eq. (18) in Ref. [Tam18030402, ] and Eq. (35) in SM of Ref. [Lam193721, ]. Under this condition, an explicit relationship between Eq. (II.1) and the Lindblad equation in the pseudomode theory can be established; see Appendix A. However, it is worth noting that Eq. (II.1) is applicable to more general situations with complex-valued $\{\eta_{k}\}$. Moreover, the Hilbert space associated with Eq. (II.1) is more compact than that of the Lindblad equation in the pseudomode theory, which is given by ${\cal H}_{\mbox{\tiny S}}\otimes{\cal H}^{\ast}_{\mbox{\tiny S}}\otimes{\cal H}_{\mbox{\tiny D}}\otimes{\cal H}_{\mbox{\tiny D}}^{\ast}$.
II.2 Exactness of bosonic DQME-SQ
To verify the exactness of the bosonic DQME-SQ theory, in the following we will give a detailed derivation of Eq. (II.1) by starting from the bosonic HEOM formalism, which is known to be an exact approach Tan20020901 . As explained in Sec. II.1, we start by expanding the hybridization correlation function of the bosonic environment as a linear combination of exponential functions,
$C(t)=\sum_{k}\eta_{k}e^{-\gamma_{k}t},$ (5)
where $\{\eta_{k}\}$ and $\{\gamma_{k}\}$ are complex numbers in general.
The exact bosonic HEOM can then be constructed as Yan16110306 ; Tan20020901 ; Wan22170901
$\dot{\rho}^{(n)}_{\bf n}$ $=\!-i\big{[}{H}_{\mbox{\tiny S}},\rho^{(n)}_{\bf n}\big{]}\!-\sum_{k}n_{k}\gamma_{k}\rho^{(n)}_{\bf n}-i\sum_{k}\!\big{[}\hat{Q},\rho^{(n+1)}_{{\bf n}_{k}^{+}}\big{]}\!$
$\quad-i\sum_{k}\!n_{k}\Big{(}\eta_{k}\hat{Q}\rho^{(n-1)}_{{\bf n}_{k}^{-}}\!-\eta_{\bar{k}}^{\ast}\rho^{(n-1)}_{{\bf n}_{k}^{-}}\hat{Q}\Big{)}.$ (6)
The dynamical variables in Eq. (II.2) are $\rho^{(n)}_{\bf n}(t)\equiv\rho^{(n)}_{n_{1}\cdots n_{K}}(t)$, where $n=n_{1}+\cdots+n_{K}$, with $n_{k}\geq 0$. In other words, each variable $\rho^{(n)}_{\bf n}(t)$ is associated with an ordered set of indexes, ${\bf n}\equiv\{n_{1}\cdots n_{K}\}$. Denote also ${\bf n}^{\pm}_{k}$ that differs from ${\bf n}$ only at the specified $k$-occupation number $n_{k}$ by ±$\pm 1$. The reduced density operator of the system is $\rho_{\mbox{\tiny S}}(t)\equiv\rho_{\bf 0}^{(0)}(t)\equiv\rho_{0\cdots 0}^{(0)}(t)$.
To proceed, we make use of two sets of complex-valued parameters, $\{\zeta_{k}\}$ and $\{\xi_{k}\}$, given below Eq. (3) and Eq. (II.1), respectively. Define dimensionless variables
$\bar{\rho}_{\bf n}^{(n)}(t)\equiv\rho_{\bf n}^{(n)}(t)/\prod_{k}\big{(}\zeta^{n_{k}}_{k}\sqrt{n_{k}!}\big{)}.$ (7)
Equation (II.2) is then recast into Li23214110
$\dot{\bar{\rho}}^{(n)}_{\bf{n}}=$ $-i[H_{\mbox{\tiny S}},\bar{\rho}^{(n)}_{{\bf n}}]-\sum_{k}{n_{k}}\gamma_{k}\bar{\rho}^{(n)}_{{\bf n}}$
$-i\sum_{k}\zeta_{k}\hat{Q}^{\times}\Big{[}\sqrt{n_{k}+1}\bar{\rho}^{(n+1)}_{{\bf n}^{+}_{k}}+\sqrt{n_{k}}\bar{\rho}^{(n-1)}_{{\bf n}^{-}_{k}}\Big{]}$
$+\sum_{k}\sqrt{n_{k}}\xi_{k}\hat{Q}^{\diamond}\bar{\rho}^{(n-1)}_{{\bf n}^{-}_{k}}.$ (8)
Here, we denote the superoperators $\hat{A}^{\times}\hat{O}\equiv\hat{A}\hat{O}-\hat{O}\hat{A}$ and $\hat{A}^{\diamond}\hat{O}\equiv\hat{A}\hat{O}+\hat{O}\hat{A}$ for any operators $\hat{A}$ and $\hat{O}$.
To derive the bosonic DQME-SQ, we take the RDT of the dissipaton-embedded system $\tilde{\rho}(t)$ as a state vector in the space $\mathcal{H}_{\mbox{\tiny S}}\otimes\mathcal{H}_{\mbox{\tiny S}}^{*}\otimes{\cal H}_{\mbox{\tiny D}}$, i.e., the reduced system space ${\cal H}_{\mbox{\tiny S}}\otimes{\cal H}^{\ast}_{\mbox{\tiny S}}$ further dilated by the dissipaton Fock space ${\cal H}_{\mbox{\tiny D}}=\otimes_{k}{\cal H}_{f_{k}}$. It reads
$\tilde{\rho}(t)\equiv\sum_{\bf n}\bar{\rho}_{\bf n}^{(n)}(t)\otimes|{\bf n}\rangle$ (9a)
with
$|{\bf n}\rangle=|n_{1}\rangle\otimes|n_{2}\rangle\otimes\cdots\otimes|n_{K}\rangle.$ (9b)
Here, $\{\bar{\rho}_{\bf n}^{(n)}(t)\}$ are defined in Eq. (7), and $|n_{k}\rangle\in\mathcal{H}_{f_{k}}$ presents the occupancy of the $k$th dissipaton level, which is generated by bosonic creation and annihilation operators $\{\hat{b}_{k}^{\pm}\}$ following the standard bosonic commutation algebra
$[\hat{b}^{-}_{k},\hat{b}^{+}_{k^{\prime}}]=\delta_{kk^{\prime}}\quad\text{and}\quad[\hat{b}^{\pm}_{k},\hat{b}^{\pm}_{k^{\prime}}]=0,$ (10)
as well as the relations,
$\hat{b}_{k}^{-}|{\bf n}\rangle=\sqrt{n_{k}}|{\bf n}^{-}_{k}\rangle,$ (11a)
$\hat{b}_{k}^{+}|{\bf n}\rangle=\sqrt{n_{k}+1}|{\bf n}^{+}_{k}\rangle.$ (11b)
Combining Eq. (II.2) and Eq. (9a), along with the relations
$\hat{b}^{-}_{k}\tilde{\rho}=\sum_{\bf n}\bar{\rho}^{(n+1)}_{{\bf n}^{+}_{k}}\otimes\hat{b}^{-}_{k}|{\bf n}^{+}_{k}\rangle=\sum_{\bf n}\sqrt{n_{k}+1}\bar{\rho}^{(n+1)}_{{\bf n}^{+}_{k}}\otimes|{\bf n}\rangle,$
$\hat{b}^{+}_{k}\tilde{\rho}=\sum_{\bf n}\bar{\rho}^{(n-1)}_{{\bf n}^{-}_{k}}\otimes\hat{b}^{+}_{k}|{\bf n}^{-}_{k}\rangle=\sum_{\bf n}\sqrt{n_{k}}\bar{\rho}^{(n-1)}_{{\bf n}^{-}_{k}}\otimes|{\bf n}\rangle,$
and
$\hat{N}_{k}\tilde{\rho}=\sum_{\bf n}\bar{\rho}^{(n)}_{{\bf n}}\otimes\hat{b}^{+}_{k}\hat{b}^{-}_{k}|{\bf n}\rangle=\sum_{\bf n}n_{k}\bar{\rho}^{(n)}_{{\bf n}}\otimes|{\bf n}\rangle,$
we then obtain the dynamical equation for the RDT $\tilde{\rho}(t)$, which is exactly the bosonic DQME-SQ of Eq. (II.1).
II.3 Universality and compactness
It is straightforward to extend the DQME-SQ of Eq. (II.1) to OQSs containing multiple degrees of freedom that are coupled to more than one environment. For instance, consider the case of $H_{\mbox{\tiny S}\mbox{\tiny E}}=\sum_{\alpha u}\hat{Q}_{u}\hat{F}_{\alpha u}$, where $\alpha$ and $u$ label the specific environment and degree of freedom, respectively. The exponential decomposition reads $C_{\alpha u}(t)\equiv\langle\hat{F}_{\alpha u}(t)\hat{F}_{\alpha u}(0)\rangle_{\mbox{\tiny E}}=\sum_{k}\eta_{\alpha uk}e^{-\gamma_{\alpha uk}t}$, and the corresponding DQME-SQ is
$\dot{\tilde{\rho}}$ $=-i[H_{\mbox{\tiny S}},\tilde{\rho}]-\sum_{\alpha uk}\gamma_{\alpha uk}\hat{b}_{\alpha uk}^{+}\hat{b}_{\alpha uk}^{-}\tilde{\rho}+\sum_{\alpha uk}\xi_{\alpha uk}\hat{b}_{\alpha uk}^{+}\{\hat{Q}_{u},\tilde{\rho}\}$
$\quad-i\sum_{\alpha uk}\zeta_{\alpha uk}(\hat{b}_{\alpha uk}^{+}+\hat{b}_{\alpha uk}^{-})[\hat{Q}_{u},\tilde{\rho}].$ (12)
Here, $\zeta_{\alpha uk}\equiv\sqrt{(\eta_{\alpha uk}+\eta^{*}_{\alpha u\bar{k}})/2}$ and $\xi_{\alpha uk}\equiv(\eta_{\alpha uk}-\eta^{*}_{\alpha u\bar{k}})/(2i\zeta_{\alpha uk})$, and $\tilde{\rho}$ is represented as a state vector in the Hilbert space $\mathcal{H}_{\mbox{\tiny S}}\otimes\mathcal{H}_{\mbox{\tiny S}}^{*}\otimes{\cal H}_{\mbox{\tiny D}}$, i.e., the system space ${\cal H}_{\mbox{\tiny S}}\otimes{\cal H}^{\ast}_{\mbox{\tiny S}}$ dilated by the dissipaton Fock space ${\cal H}_{\mbox{\tiny D}}=\otimes_{\alpha uk}{\cal H}_{f_{\alpha uk}}$.
As evidenced in Appendix A, the bosonic DQME-SQ fully captures the non-Markovian open quantum dynamics within a smaller Hilbert space compared to the pseudomode theory, without any approximation.
III Fermionc DQME-SQ theory
III.1 Construction of fermionic DQME-SQ